Wright Corporation developed the following standard costs for one of the products of their products:

Standard Cost per Unit

Materials: 4ft x $14.25 per foot $57

Direct labor: 8 hours x $10 per hour $80

Variable Overhead: 8 direct labor hours x $8 per hour $64

Fixed Overhead: 8 direct labor hours x $12 per hour $96

Total standard cost per unit $297

The following information is available regarding the company's operations for the period:

Units Produced 11,000

Materials Purchased 52,000 ft @ $13.95 per foot

Materials used: 40,000 ft

Direct labor: 84,000 hours costing $840,000

Manufacturing overhead incurred:

Variable $756,000

Fixed $1,000,000

Budgeted fixed manufacturing overhead for the period is $960,000 and the standard fixed overhead rate is base expected capacity of 80,000 direct labor hours.

Required:

A. Calculate the materials price variance

B. Calculate the materials usage variance

C. Calculate the direct labor rate variance

D. Calculate the direct labor efficiency variance

E. Calculate the variable manufacturing overhead spending variance

F. Calculate the variable manufacturing overhead efficiency variance

G. Calculate the fixed manufacturing overhead spending variance

H. Calculate the fixed manufacturing overhead volume variance
Answer #1

A. Materials price variance = AQP x (AP - SP) = 52000 x ($13.95 - $14.25) = $15600 Favorable

B. Materials usage variance = SP x (AQ - SQ) = $14.25 x (40000 - 44000) = $57000 Favorable

SQ = 11000 units x 4ft = 44000

C. Direct labor rate variance = AH x (AR - SR) = 84000 x ($10 - $10) = $0

AR = $840000/84000 = $10

D. Direct labor efficiency variance = SR x (AH - SH) = $10 x (84000 - 88000) = $40000 Favorable

SH = 11000 units x 8 hours = 88000

Per HOMEWORKLIB RULES the first 4 parts have been answered. Please post the remaining parts separately. Thank you.